How many grams of copper are required to react with 45.0 grams of silver nitrate? Cu+ 2 AgNO3= Cu(NO3)2 + 2 Ag


1
Expert's answer
2021-01-11T03:49:12-0500

Solution:

The balanced chemical equation:

Cu + 2AgNO3= Cu(NO3)2 + 2Ag

According to the equation: n(Cu) = n(AgNO3)/2


Moles of AgNO3 = Mass of AgNO3 / Molar mass of AgNO3

The molar mass of AgNO3 is 169.87 g mol-1.

n(AgNO3) = Moles of AgNO3 = 45.0 g / 169.87 g mol-1 = 0.2649 mol


n(Cu) = n(AgNO3)/2 = 0.2649 mol / 2 = 0.13245 mol


Moles of Cu = Mass of Cu / Molar mass of Cu

The molar mass of Cu is 63.546 g mol-1.

Hence,

Mass of Cu = Moles of Cu × Molar mass of Cu = 0.13245 mol × 63.546 g mol-1 = 8.41667 g = 8.42 g

Mass of Cu is 8.42 grams.


Answer: 8.42 grams of copper (Cu) are required.